OpenStudy (anonymous):

so confused..."write the equation of a line, in standard form, passing through the point A(-2,7) and parallel to the line whose equation is y=2x-3" ..help please??

OpenStudy (anonymous):

parallel lines have the same slope the slope of \(y=mx+b\) is \(m\) and therefore the slope of \(y=2x-3\) is 2

OpenStudy (anonymous):

now you use the point - slope formula \[y=y_1=m(x-x_1)\] with \(m=2, x_1=-2, y_1=7\) make the real number replacements and see what you get

OpenStudy (anonymous):

typo there, should have been \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

you get \[y-7=2(x-(-2))\] or \[y-7=2(x+2)\]

OpenStudy (anonymous):

thank you